OpenStudy (lovelyharmonics):

Write the expression in standard form. six divided by quantity four minus thirteen i.

OpenStudy (anonymous):

\[\frac{6}{4-13i}\]?

OpenStudy (lovelyharmonics):

yes c:

OpenStudy (anonymous):

multiply top and bottom by the conjugate of the denominator the conjugate of \(a+bi\) is \(a-bi\) and this works because \((a+bi)(a-bi)=a^2+b^2\) a real number

OpenStudy (anonymous):

in your case you have \(4-13i\) and the conjugate is \(4+13i\) start with \[\frac{6}{4-13i}\times \frac{4+13i}{4+13i}=\frac{6(4+13i)}{4^2+13^2}\]

OpenStudy (anonymous):

you probably want to multiply out in the top and compute the bottom also

OpenStudy (lovelyharmonics):

24+78i?

OpenStudy (anonymous):

that is the numerator also compute the denominator, then break apart to write in standard form

OpenStudy (anonymous):

i get \(185\) for \(4^2+13^2\) so it is \[\frac{24+78i}{185}\] standard form is \(a+bi\) so final answer is \[\frac{24}{185}+\frac{78}{185}i\]